Given that tanθ = 1/4√3 show that sin(120+θ)=11/14
Answer by CharlesG2(834)

Given that tanθ = 1/4√3 show that sin(120+θ)=11/14

tan θ = 1/(4√3)
θ = 8.213211 degrees to 6 places
sin (120 + 8.213211) = sin (128.213211)
sin (128.213211) = 0.785714 to 6 places
11/14 = 0.785714 to 6 places
sin(120+θ)=11/14